OpenStudy (anonymous):

sarah had 7/8 pound of turkey she used 2/3 pound for sandwitches how many pounds of turkey does sarah have left

OpenStudy (calculusxy):

\[\frac{ 7 }{ 8 } - \frac{ 2 }{ 3 } = \frac{ 21 }{ 24 } - \frac{ 16 }{ 24 } = \frac{ 5 }{ 24 }\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can you help whith 1 more